OpenStudy (anonymous):

How many cones of radius 1 m and height 1 m have total volume equal to the volume of a sphere with radius 1 m?

OpenStudy (raden):

first, you have to find out the volume of them respectively use the formula : Volume of cone = 1/3 pi r^2 h Volume of sphere = 4/3 pi r^3